OpenStudy (anonymous):

$8,000 is invested in two different accounts earning 3% and 4% interest. If the two accounts earn a total of $295 in interest, how much money was invested at 3%? (idk wat the answer is)

OpenStudy (anonymous):

Let A be the amount invested at 3%, B the amount at 4%. Total value: 1) A+B=8000 Interest: 0.03A+0.04B=295 2) 3A+4B=29500 Multiply eq. 1 by -3 and add to eq. 2, -3A-3B+3A+4B=-24000+29500 B=5500 Then from eq. 1, A+5500=8000 A=8000-5500 A=2500

OpenStudy (anonymous):

thx @cutepochacco